# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ III.2 Symmetry and topological principles to probe monopole superconducting order
Symmetry principles have long been crucial in the design of Josephson experiments to reveal the pairing symmetry of unconventional superconductors [36, 16, 20, 21]. We now propose symmetry and topological principles to guide the design of a set of Josephson junctions which together can identify monopole superconducting order. This approach complements the discussion in Sec. III.1, which was mainly based on microscopic derivations of Josephson current.
@@
Monopole pairing order with pair monopole charge $q_{p}$ generally takes the form of a superposition of monopole harmonic functions $\mathcal{Y}_{q_{p},l,l_{\hat{\mathbf{n}},\mathrm{glob}}}$ for different integer values of $l\geq|q_{p}|$ and $|l_{\hat{\mathbf{n}}}|\leq l$. For $q_{p}\neq 0$, the monopole harmonic functions exhibit nodes pinned to the rotation quantization axis, $\hat{\mathbf{n}}$, which we often take as $\hat{z}$ without loss of generality. Furthermore, as discussed in Sec. III.1, local angular momenta differ from global angular momentum by the pair monopole charge,
@@
(I) Globally, the monopole superconducting order remains a representation of the rotation group but resides in a topological sector characterized by its pair monopole charge $q_{p}$. For two superconductors that transform identically under a given symmetry operation, which is also a symmetry of the junction, the Josephson energy phase relationship is invariant under this symmetry transformation. Hence, if rotation $R_{z}$ with respect to the $z$ axis is a symmetry of the junction, then the first order Josephson coupling is allowed only if the superconducting orders at both sides of the junction transform in the same way under $R_{z}$, i.e. they both form an irrep of $R_{z}$ labeled by the same $l_{z}$. Monopole superconducting order given by, for example, $\mathcal{Y}_{q_{p};l,l_{z,\mathrm{glob}}}$ transforms under rotation $R_{z}$ according to its globally conserved angular momentum, $l_{z,\mathrm{glob}}$. As such, to extract $l_{z,\mathrm{glob}}$, we consider a junction which is oriented along a rotational $z$ axis, as shown in Fig. 1 (a). By placing a monopole superconductor in junction with another superconductor with known pairing symmetry, one can extract the global angular momentum $l_{z,\mathrm{glob}}$ based on the existence of first order $2\pi$-periodic Josephson current.
@@
(II)  Locally in momentum space, the symmetry of monopole superconductor pairing order can be described by regular spherical harmonic symmetry. In the local representation, the pairing phase possesses nontrivial winding (see Appendix C), corresponding to chiral pairing order characterized by angular momentum $l_{z,\mathrm{loc}}$. Between the gap nodes pinned to the rotational axis, the local chirality changes by twice the pair monopole charge, as shown in the model in Sec. III.1. To probe $l_{z,\mathrm{loc}}$, we consider the design shown in Fig. 1 (b), in which two identical monopole superconductors form a Josephson junction oriented normal to the $z$ axis. For given conserved momentum $k_{z}$ along the rotational axis, the junction can be regarded effectively as a junction between two identical two-dimensional superconductors whose pairing orders are characterized by local angular momentum $l_{z,\mathrm{loc}}$. Consequently, the total Josephson current of the proposed junction is a superposition of contributions from these planar junctions, which can give signatures, for example the periodicity and phase shift of the current phase relation, that can indicate the local pairing angular momentum $l_{z,\mathrm{loc}}$.
@@
In the following two sections, we analyze a set of Josephson junctions guided by the above design principles to probe the pairing symmetry of a prototypical monopole superconductor arising in a magnetic doped Weyl semimetal with inter-Fermi surface pairing in Eq. (18). In Sec. IV and Sec. V, we propose junctions to probe the global angular momentum $l_{z,\mathrm{glob}}$ and local angular momentum $l_{z,\mathrm{loc}}$ respectively. Together, these junctions can be used to extract the monopole charge, $|q_{p}|=|l_{z,\mathrm{glob}}-l_{z,\mathrm{loc}}|$, and provide a foundation to guide efforts to experimentally uncover the monopole superconducting order.
@@ IV Josephson junctions to probe $l_{z,\mathrm{glob}}$ of the monopole superconducting order
Following the symmetry principles presented in Sec. III.2, we design Josephson junctions to extract the global angular momentum of monopole superconducting order, $l_{z,\mathrm{glob}}$. We first examine a Josephson junction oriented along the $z$ direction between a monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ and a conventional $s$-wave superconductor, demonstrating the nonvanishing first order Josephson current. After, in Sec. IV.2, we study the Josephson coupling for the monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ in the same topological sector but with different global angular momentum, and in Sec. IV.3, monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ in a different topological sector of $q_{p}=3$.
IV.1 First order Josephson current between $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor
@@ IV.1.1 Method 1: Linear Response Theory
We study Josephson effects between an $s$-wave superconductor and monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ in a junction along the $z$ direction, shown schematically in Fig. 2 (a). Following the symmetry principles, as discussed in Sec. III.2, the first order Josephson coupling in this system is allowed due to the pairing order at two sides of the junction being characterized by the same global angular momentum $l_{z}=0$. To show this, we consider the continuum model of the monopole superconductor introduced in Eq. (18), for which the effective pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ in Eq. (24) is invariant under rotation $R_{z}$, having global angular momentum $l_{z}=0$. Following the derivation based on linear response theory in Sec. III.1, we demonstrate nonzero first order Josephson coupling and the resulting critical current.
@@
where $\Delta_{L,0}$ is the pairing amplitude of the $s$-wave superconductor, and $\Delta_{R,0}$ is the amplitude of the inter-Fermi surface pairing in the monopole superconductor. We model the tunneling amplitude through the junction barrier as a spin-independent delta function that preserves in-plane momenta with $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}}$ and $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=0$. Here, $T_{0}$ is a constant, and $\mathbf{k}_{\parallel}=\mathbf{k}_{\parallel}^{\prime}=\tilde{\mathbf{k}}_{\parallel}$ is the wave vector in the $k_{x}k_{y}$ plane, which is conserved at the barrier. Already, the nonvanishing first order Josephson coupling can be seen from the form factor $\mathfrak{F}_{s-\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$, which encodes the symmetry of the junction as well as the pairing orders and itself is invariant under rotation $R_{z}$ and parity.
@@
describes the contribution to the Josephson current for conserved transverse momentum, ${k}_{\parallel}$. Above, $A$ is the surface area of the interface, and $\ell_{L}$ and $\ell_{R}$ are the system lengths along $z$ direction at the left and right sides respectively. We consider the Josephson current at zero temperature, for which $w(E_{L,\mathbf{k}^{\prime}},E_{R,\tilde{\mathbf{k}}}^{(bp)};\beta)=[2E_{L,\mathbf{k}^{\prime}}E_{R,\tilde{\mathbf{k}}}^{(bp)}(E_{L,\mathbf{k}^{\prime}}^{2}-E_{R,\tilde{\mathbf{k}}}^{(bp)2})]^{-1}$ and is positive-definite. The dispersions of the BdG quasiparticles for the $s$-wave superconductor and monopole superconductor are given by $E_{L,\mathbf{k}^{\prime}}=\sqrt{\xi_{L,\mathbf{k}^{\prime}}^{2}+|\Delta_{L,0}|^{2}}$ and $E_{R,\tilde{\mathbf{k}}}^{(bp)}=\sqrt{\xi_{R,\tilde{\mathbf{k}}}^{(bp)2}+\left|\Delta_{R,0}\right|^{2}{\tilde{k}^{2}_{\parallel}}/{\tilde{k}^{2}}}$ respectively. Here, $\xi_{L,\mathbf{k}^{\prime}}=\hbar^{2}({k^{\prime}_{z}}^{2}+{k^{\prime}}_{\parallel}^{2})/2m-\mu_{L}$ and $\xi^{(bp)}_{R,\tilde{\mathbf{k}}}=\hbar v_{F}\sqrt{\tilde{k}_{z}^{2}+\tilde{k}_{\parallel}^{2}}-\mu_{R}$ are the respective band dispersions. The Fermi surfaces of $\xi_{L,\mathbf{k}^{\prime}}$ and $\xi_{R,\tilde{\mathbf{k}}}^{(bp)}$ are shown projected to the $k_{x}k_{y}$ plane in Fig. 2 (b).
@@
Figure 2: (a) Schematic of the Josephson junction oriented along $z$ direction between an $s$-wave superconductor and a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. (b) Bulk Fermi surface projections on the $k_{x}k_{y}$ plane for $s$-wave superconductor (hatched) and monopole superconductor (solid) at two sides of the junction in (a). (c) Relative magnitude of the contributions to the first order Josephson current phase relation for different transverse momentum, $k_{\parallel}$, corresponding to Eq. (32). (d) First order Josephson current between the monopole superconductor and $s$-wave superconductor at zero temperature and zero bias. Here, $v_{F}=0.5$, $\mu_{R}=0.1$, $\tilde{k}_{F}=0.2$, and $\mu_{L}=0.8$. $\Delta_{L,0}=\Delta_{R,0}=\Delta_{0}=0.005$, $I_{c}=0.62\Delta_{0}e/\hbar$, and $N_{k_{\parallel}}$ is the number of transverse momentum cuts in the calculation. $T_{0}$, $A$, $\ell_{L}$, $\ell_{R}$, and $m$ are set to unity.
@@
In Fig. 2 (c), we show $J(k_{\parallel})$, the contributions to the first order Josephson current, for different conserved transverse momentum $k_{\parallel}$. States near the equator of the Fermi surface of the doped Weyl semimetal, with $k_{\parallel}\approx\tilde{k}_{F}$, contribute the most to the Josephson tunneling, where the magnitude of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ is at its maximum. As $k_{\parallel}$ approaches zero, the contributions to the Josephson current are suppressed as the magnitude of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ reaches its minimum near its gap nodes along the $k_{z}$ axis. For transverse momentum away from the Weyl Fermi surfaces, $k_{\parallel}>\tilde{k}_{F}+{\Delta_{R,0}}/{v_{F}}$, its contribution to the Josephson current vanishes due to vanishing low energy density of states in $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor.
@@
Furthermore, in Fig. 2 (d), we show the total first order Josephson current phase relation given in Eq. (31) numerically. It exhibits $2\pi$ periodicity and nonvanishing critical current $I_{c}=0.62\Delta_{0}e/\hbar$. In contrast, for a Josephson junction in the same geometry between an $s$-wave and a chiral $p_{x}\pm ip_{y}$-wave superconductor, the first order Josephson current vanishes when there is no spin-orbit coupling in the bulk nor at the interface. Hence, though locally $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ exhibits chiral $p_{x}\pm ip_{y}$ pairing symmetry, the Josephson current in the above junction serves as an effective probe of the global angular momentum of the pairing order $l_{z,\mathrm{glob}}=0$. We discuss later the case when there exist spin-orbit interactions in Sec. VI.

### v2

@@ III.2 Symmetry and topological principles to probe monopole superconducting order
Symmetry principles have long been crucial in the design of Josephson experiments to reveal the pairing symmetry of unconventional superconductors Geshkenbein and Larkin (1986); Geshkenbein et al. (1987); Van Harlingen (1995); Tsuei and Kirtley (2000). We now propose symmetry and topological principles to guide the design of a set of Josephson junctions which together can identify monopole superconducting order. This approach complements the discussion in Sec. III.1, which was mainly based on microscopic derivations of the Josephson current.
@@
Monopole pairing order with pair monopole charge $q_{p}$ generally takes the form of a superposition of monopole harmonic functions $\mathcal{Y}_{q_{p},l,l_{\hat{\mathbf{n}},\mathrm{glob}}}$ for different integer values of $l\geq|q_{p}|$ and $|l_{\hat{\mathbf{n}}}|\leq l$. For $q_{p}\neq 0$, the monopole harmonic functions exhibit nodes pinned to the rotation quantization axis, $\hat{\mathbf{n}}$, which we often take as $\hat{z}$ without loss of generality. Furthermore, as discussed in Sec. III.1, local angular momenta differ from the global angular momentum by the pair monopole charge,
@@
(i) Globally, the monopole superconducting order remains a representation of the rotation group but resides in a topological sector characterized by its pair monopole charge $q_{p}$. For two superconductors that transform identically under a given symmetry operation, which is also a symmetry of the junction, the Josephson energy phase relationship is invariant under this symmetry transformation. Hence, if rotation $R_{z}$ with respect to the $z$ axis is a symmetry of the junction, then the first-order Josephson coupling is allowed only if the superconducting orders at both sides of the junction transform in the same way under $R_{z}$, i.e. they both form an irrep of $R_{z}$ labeled by the same $l_{z}$. Monopole superconducting order given by, for example, $\mathcal{Y}_{q_{p};l,l_{z,\mathrm{glob}}}$ transforms under rotation $R_{z}$ according to its globally conserved angular momentum, $l_{z,\mathrm{glob}}$. As such, to extract $l_{z,\mathrm{glob}}$, we consider a junction which is oriented along a rotational $z$ axis, as shown in Fig. 1(a). By placing a monopole superconductor in junction with another superconductor with known pairing symmetry, one can extract the global angular momentum $l_{z,\mathrm{glob}}$ based on the existence of first-order $2\pi$-periodic Josephson current.
@@
(ii)  Locally in momentum space, the symmetry of monopole superconductor pairing order can be described by regular spherical harmonic symmetry. In the local representation, the pairing phase possesses nontrivial winding (see Appendix C), corresponding to chiral pairing order characterized by angular momentum $l_{z,\mathrm{loc}}$. Between the gap nodes pinned to the rotational axis, the local chirality changes by twice the pair monopole charge, as shown in the model in Sec. III.1. To probe $l_{z,\mathrm{loc}}$, we consider the design shown in Fig. 1(b), where two identical monopole superconductors form a Josephson junction oriented normal to the $z$ axis. For given conserved momentum $k_{z}$ along the rotational axis, the junction can be regarded effectively as a junction between two identical two-dimensional superconductors whose pairing orders are characterized by local angular momentum $l_{z,\mathrm{loc}}$. Consequently, the total Josephson current of the proposed junction is a superposition of contributions from these planar junctions, which can give signatures, for example the periodicity and phase shift of the current phase relation, that can indicate the local pairing angular momentum $l_{z,\mathrm{loc}}$.
@@
In the following two sections, we analyze a set of Josephson junctions guided by the above design principles to probe the pairing symmetry of a prototypical monopole superconductor arising in a magnetic doped Weyl semimetal with inter-Fermi surface pairing in Eq. (18). In Secs. IV and V, we propose junctions to probe the global angular momentum $l_{z,\mathrm{glob}}$ and local angular momentum $l_{z,\mathrm{loc}}$ respectively. Together, these junctions can be used to extract the monopole charge, $|q_{p}|=|l_{z,\mathrm{glob}}-l_{z,\mathrm{loc}}|$, and provide a foundation to guide efforts to experimentally uncover the monopole superconducting order.
@@ IV Josephson junctions to probe $l_{z,\mathrm{glob}}$ of the monopole superconducting order
Following the symmetry principles presented in Sec. III.2, we design Josephson junctions to extract the global angular momentum of monopole superconducting order, $l_{z,\mathrm{glob}}$. We first examine a Josephson junction oriented along the $z$ direction between a monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ and a conventional $s$-wave superconductor, demonstrating the nonvanishing first-order Josephson current. After, in Sec. IV.2, we study the Josephson coupling for the monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ in the same topological sector but with different global angular momentum, and in Sec. IV.3, monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ in a different topological sector of $q_{p}=3$.
IV.1 First-order Josephson current between $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor
@@ IV.1.1 Method 1: linear response theory
We study Josephson effects between an $s$-wave superconductor and a monopole superconductor with pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ in a junction along the $z$ direction, shown schematically in Fig. 2(a). Following the symmetry principles, as discussed in Sec. III.2, the first-order Josephson coupling in this system is allowed due to the pairing order at two sides of the junction being characterized by the same global angular momentum $l_{z}=0$. To show this, we consider the continuum model of the monopole superconductor introduced in Eq. (18), for which the effective pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ in Eq. (24) is invariant under rotation $R_{z}$, having global angular momentum $l_{z}=0$. Following the derivation based on linear response theory in Sec. III.1, we demonstrate nonzero first-order Josephson coupling and the resulting critical current.
@@
where $\Delta_{L,0}$ is the pairing amplitude of the $s$-wave superconductor, and $\Delta_{R,0}$ is the amplitude of the inter-Fermi surface pairing in the monopole superconductor. We model the tunneling amplitude through the junction barrier as a spin-independent delta function that preserves in-plane momenta with $T_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}\delta_{\mathbf{k}_{\parallel},\mathbf{k}^{\prime}_{\parallel}}$ and $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})=0$. Here, $T_{0}$ is a constant, and $\mathbf{k}_{\parallel}=\mathbf{k}_{\parallel}^{\prime}=\tilde{\mathbf{k}}_{\parallel}$ is the wave vector in the $k_{x}k_{y}$ plane, which is conserved at the barrier. Already, the nonvanishing first-order Josephson coupling can be seen from the form factor $\mathfrak{F}_{s-\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$, which encodes the symmetry of the junction as well as the pairing orders and itself is invariant under rotation $R_{z}$ and parity.
@@
describes the contribution to the Josephson current for conserved transverse momentum, ${k}_{\parallel}$. Here, $A$ is the surface area of the interface, and $\ell_{L}$ and $\ell_{R}$ are the system lengths along $z$ direction at the left and right sides respectively. We consider the Josephson current at zero temperature, for which $w(E_{L,\mathbf{k}^{\prime}},E_{R,\tilde{\mathbf{k}}}^{(bp)};\beta)=[2E_{L,\mathbf{k}^{\prime}}E_{R,\tilde{\mathbf{k}}}^{(bp)}(E_{L,\mathbf{k}^{\prime}}^{2}-E_{R,\tilde{\mathbf{k}}}^{(bp)2})]^{-1}$ and is positive-definite. The dispersions of the BdG quasiparticles for the $s$-wave superconductor and monopole superconductor are given by $E_{L,\mathbf{k}^{\prime}}=\sqrt{\xi_{L,\mathbf{k}^{\prime}}^{2}+|\Delta_{L,0}|^{2}}$ and $E_{R,\tilde{\mathbf{k}}}^{(bp)}=\sqrt{\xi_{R,\tilde{\mathbf{k}}}^{(bp)2}+\left|\Delta_{R,0}\right|^{2}{\tilde{k}^{2}_{\parallel}}/{\tilde{k}^{2}}}$ respectively. Here, $\xi_{L,\mathbf{k}^{\prime}}=\hbar^{2}({k^{\prime}_{z}}^{2}+{k^{\prime}}_{\parallel}^{2})/2m-\mu_{L}$ and $\xi^{(bp)}_{R,\tilde{\mathbf{k}}}=\hbar v_{F}\sqrt{\tilde{k}_{z}^{2}+\tilde{k}_{\parallel}^{2}}-\mu_{R}$ are the respective band dispersions. The Fermi surfaces of $\xi_{L,\mathbf{k}^{\prime}}$ and $\xi_{R,\tilde{\mathbf{k}}}^{(bp)}$ are shown projected to the $k_{x}k_{y}$ plane in Fig. 2(b).
@@
Figure 2: (a) Schematic of the Josephson junction oriented along $z$ direction between an $s$-wave superconductor and a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. (b) Bulk Fermi surface projections on the $k_{x}k_{y}$ plane for $s$-wave superconductor (hatched) and monopole superconductor (solid) at two sides of the junction in (a). (c) Relative magnitude of the contributions to the first-order Josephson current phase relation for different transverse momentum, $k_{\parallel}$, corresponding to Eq. (32). (d) first-order Josephson current between the monopole superconductor and $s$-wave superconductor at zero temperature and zero bias. Here, $v_{F}=0.5$, $\mu_{R}=0.1$, $\tilde{k}_{F}=0.2$, and $\mu_{L}=0.8$. $\Delta_{L,0}=\Delta_{R,0}=\Delta_{0}=0.005$, $I_{c}=0.62\Delta_{0}e/\hbar$, and $N_{k_{\parallel}}$ is the number of transverse momentum cuts in the calculation. $T_{0}$, $A$, $\ell_{L}$, $\ell_{R}$, and $m$ are set to unity.
@@
In Fig. 2(c), we show $J(k_{\parallel})$, the contributions to the first-order Josephson current, for different conserved transverse momentum $k_{\parallel}$. States near the equator of the Fermi surface of the doped Weyl semimetal, with $k_{\parallel}\approx\tilde{k}_{F}$, contribute the most to the Josephson tunneling, where the magnitude of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ is at its maximum. As $k_{\parallel}$ approaches zero, the contributions to the Josephson current are suppressed as the magnitude of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ reaches its minimum near its gap nodes along the $k_{z}$ axis. For transverse momentum away from the Weyl Fermi surfaces, $k_{\parallel}>\tilde{k}_{F}+{\Delta_{R,0}}/{v_{F}}$, its contribution to the Josephson current vanishes due to vanishing low-energy density of states in $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor.
@@
Furthermore, in Fig. 2(d), we show the total first-order Josephson current phase relation given in Eq. (31) numerically. It exhibits $2\pi$ periodicity and nonvanishing critical current $I_{c}=0.62\Delta_{0}e/\hbar$. In contrast, for a Josephson junction in the same geometry between an $s$-wave and a chiral $p_{x}\pm ip_{y}$-wave superconductor, the first-order Josephson current vanishes when there is no spin-orbit coupling in the bulk nor at the interface. Hence, though locally $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ exhibits chiral $p_{x}\pm ip_{y}$ pairing symmetry, the Josephson current in the above junction serves as an effective probe of the global angular momentum of the pairing order $l_{z,\mathrm{glob}}=0$. We discuss later the case when there exist spin-orbit interactions in Sec. VII.